OpenStudy (anonymous):

Cos^2theta - Sin^2theta=?

OpenStudy (anonymous):

Trigonometric Identities

OpenStudy (sidsiddhartha):

start with \[\cos(a+b)=cosa*cosb-sina*sinb\]

OpenStudy (anonymous):

Welcome to Openstudy.

OpenStudy (sidsiddhartha):

are u familiar with this formula?

OpenStudy (anonymous):

I only have 5 possible answers also! \[\left( \cos \theta -\sin \theta \right)^{2}\]

OpenStudy (sidsiddhartha):

\[\cos(a+b)=cosa∗cosb−sina∗sinb\] yeah in this indentity if u repalce b=a then u will get \[\cos(a+a)=cosa*cosa-sina*sina \] \[\cos(2a)=\cos^2(a)-\sin^2(a)\] so \[\cos^2(\theta)-\sin^2(\theta)=\cos(2\theta)\] ok????